OpenStudy (yacoub1993):

Solve the equation for y. Then graph the equation. 5x – 4y = 12 A. y = -5/4x - 3 B. y = 4/5x + 3 C. y = 5/4x - 3 D. y = -4/5x + 3

OpenStudy (anonymous):

take the terms of x on right side first

OpenStudy (anonymous):

do it asap

OpenStudy (yacoub1993):

-4y=12-5x

OpenStudy (anonymous):

divide both sides by -4

OpenStudy (anonymous):

tell me wat u got

OpenStudy (yacoub1993):

i got y=\[\frac{ 12-5x }{-4 }\]

OpenStudy (anonymous):

check wid the options u hav mentioned asap

OpenStudy (anonymous):

asap

OpenStudy (yacoub1993):

\[\frac{ -4y }{-4 }=\frac{ 12-5x }{-4 }\]

OpenStudy (anonymous):

perform the division quickly

OpenStudy (yacoub1993):

didnt get from here

OpenStudy (yacoub1993):

@manishsatywali i didt get it from here Any Idea

OpenStudy (anonymous):